OpenStudy (anonymous):

Find the specified vector or scalar. Show your work. u = -4i + 1j and v = 4i + 1j; Find .||u+v||

OpenStudy (anonymous):

@freckles

OpenStudy (kainui):

Find what?

OpenStudy (anonymous):

idk let me ask my teacher thats how the question was given to me

OpenStudy (anonymous):

||u+v||

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

(ai+bj)+(ci+dj) =(a+c)i+(b+d)j then to find the magnitude of that sum you do \[\sqrt{(a+c)^2+(b+d)^2}\]

OpenStudy (anonymous):

ok this is my first time doing this soo

OpenStudy (anonymous):

i dont know what to do ?

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@skullpatrol

OpenStudy (freckles):

add the i-components

OpenStudy (freckles):

-4+4=?

OpenStudy (freckles):

then add the j-components 1+1=?

OpenStudy (freckles):

\[u+v=(-4+4)i+(1+1)j\]

OpenStudy (anonymous):

1+1= 2 and -4+40

OpenStudy (anonymous):

sorry -4+4=0

OpenStudy (freckles):

ok so you saying u+v=0i+2j or just u+v=2j

OpenStudy (freckles):

now how far is that from the origin ?

OpenStudy (freckles):